OpenStudy (anonymous):

If cos Θ = -4/7, what are the values of sin Θ and tan Θ?

OpenStudy (anonymous):

I closed this by accident but I still need help @wolf1728 =c if you can

OpenStudy (wolf1728):

If the cosine of of Θ = -4/7 or -0.5714285714 Then Θ = arc cosine (-0.5714285714) = 124.85 Degrees so sin Θ or 124.85 degrees = 0.82065 and tan Θ or 124.85 degrees = -1.4361 hope that helps

OpenStudy (anonymous):

Would that be sin = +/- sqrt33/7 and tan = +/- sqrt33/4?

OpenStudy (wolf1728):

or cosine Θ = -4/7 cosine Θ = adjacent / hypotenuse If adjacent = -4 and hypotenuse = 7 then opposite² = hyp² - adj² opposite² = 49 -16 opposite² = 33 opposite = 5.7445626465 sine = opp / hyp and tan = opp / adj so sine Θ = 5.7445626465 / 7 and tan Θ = 5.7445626465 / -4

OpenStudy (wolf1728):

sin Θ = sqrt(33) / 7 tan Θ = sqrt(33) /-4

OpenStudy (anonymous):

Thank you! Could you help me with one last one?: http://openstudy.com/study#/updates/537cfcc8e4b0baf62d52d909